# CodeSSM: Towards State Space Models for Code Understanding

### v2

@@ 4.3 Token classification task (Type inference)
Column 5 of Table 2 shows the results on the type inference benchmark. CodeSSM (across all context lengths) matches or exceeds CodeBERT. However, both RoCoder and BertCoder outperform CodeSSM. We provide reasons for poor performance of CodeSSM on type inference in Section 5.5 and Appendix G.
@@ 4.4 Clone Detection
Column 4 of Table 2) shows the F1 on clone detection task. The CodeSSM-sc with a pretraining context length of 256 outperforms both CodeBERT, BertCoder, and RoCoder in terms of F1. Under a similar training setup (2048 context on CSN), removing positional embedding does have a minimal impact on performance, but with the advantage that the maximum input length need not be fixed during pretraining. The decrease in performance after introducing dropout shows the model does not overfit despite the dataset being unbalanced.
@@ 5.1 How sample efficient is CodeSSM?
Figure 2 shows the MLM training accuracy trajectories of CodeSSM and BertCoder, both trained with a maximum context of 2048. CodeSSM demonstrates rapid initial learning, achieving more than 50% accuracy in only 3,000 training steps (96,000 training samples). BertCoder, on the other hand, exhibits better long-term learning characteristics, outperforming CodeSSM by 4% points after six training epochs (see Appendix E). However, CodeSSM outperforms transformers in all downstream tasks except type inference, as discussed in Section 4. The performance on downstream tasks shows that the code modeling capability observed during pretraining transfers well to multiple downstream tasks.
@@
Our analyses demonstrate that SSMs represent a sample-efficient alternative to transformers for code language modeling and understanding tasks. We attribute the sample efficiency of CodeSSM to the distinct inductive bias of SSM models compared to transformer architectures Wang et al. (2023a). We also present some (informal) theoretical basis for CodeSSM’s performance and the magnitude and phase spectrum of Fourier Transform of SSM kernel in Appendix F.
@@ 5.2 How well does CodeSSM extrapolate to longer contexts?
We studied the transformer models and CodeSSMs for positional embedding and different pretraining context lengths to investigate their length extrapolation capabilities.
@@
Out of two retrieval tasks, RoCoder performs slightly better on NLCodeSearch while CodeSSM outperforms it on SQA. The context length is 512 for the former and 1024 for the latter. Thus, CodeSSM is better at handling unseen input lengths as the context for the downstream task increases. Since SSMs have linear complexity with respect to input length, they are advantageous in terms of both computational efficiency and performance. Moreover, with a longer pretraining context, CodeSSM outperforms RoCoder on NLCodeSearch as well. Therefore, with more computational resources, CodeSSM can be efficiently trained with longer contexts to enhance the performance on retrieval tasks. However, it’s important to note that longer pretraining contexts do not benefit all downstream tasks; in fact, performance slightly decreases on vulnerability detection and clone detection.
@@ 5.3 Can SSMs be replaced by DFT?
We ask this question because Lee-Thorp et al. (2022) showed that self-attention can be replaced with DFT in transformers with minimal reduction on NLP tasks. We also observed that in all tasks RoCoder performed significantly better than BertCoder. Hua et al. (2025b) showed that RoPE also performs token-level (non-uniform) DFT. Additionally, SSMs also implicitly perform DFT of input and kernel during convolution operation.
To answer this question, we performed an ablation study by replacing the SSM transformation in CodeSSM with DFT and creating a unidirectional variant of CodeSSM. The results are presented in Figure 3 (and in Table 7). We observe that both variants perform significantly worse compared to CodeSSM. However, it is important to note that CodeF performs better than CodeSSM-uni, i.e., the advantages of SSM over DFT on code modeling are only realized with bidirectional SSMs.
@@ 5.4 Is CodeSSM robust to overfitting on small datasets?
We study the robustness of the CodeSSM model in two ways. First, we study the impact of adding dropout to the CodeSSM layer. Across all tasks, including the complexity prediction task which has only 3.6k samples, the performance of the model reduces when dropout is added. The performance degradation suggests that CodeSSM does not overfit during training regardless of the size of the finetuning dataset. Second, we evaluated the model on the adversarial NLCodeSearch task, by which informative tokens such as function and variable names have been replaced with noninformative generic tokens. The CodeSSM model performs very well in this task even with minimal training.
@@ 5.5 When does CodeSSM fail and why?
As discussed in Section 5.1, transformers can outperform SSMs as the number of samples increases. This phenomenon can hinder the performance of SSMs on some tasks such as Type Inference. This benchmark consists of over 9M type annotations, which is significantly larger than the samples of other tasks. The total number of predictions (hence, feedback to the model through loss) is even larger since there is one prediction for every token. However, when the number of samples is large and the context length of finetuning datasets is long, SSMs can surpass transformers in such scenarios (as seen in the SQA benchmark).
SSMs can effectively capture long-range dependencies but do not model local dependencies (Zuo et al., 2024) very well. The type inference benchmark requires the model to understand both short and long-range dependencies. For example, in the code sample x = y+1; z = func(x), the type of x can be inferred from the local context (variable y), while the type of z can only be inferred from the return type of the function func, which requires capturing long-range dependencies. We also observe that this problem becomes more challenging when the local context is complicated. We provide a qualitative example of CodeSSM’s failure in the task of type inference in Appendix G. We can potentially improve CodeSSM’s performance on this task by making the step size of SSMs (see Appendix A) data dependent Qi et al. (2024) or with multiple SSM kernels (Appendix G), which can improve its understanding of short-range dependencies.
@@ 6 Related Work
Transformer models have been used in code intelligence tasks Feng et al. (2020); Guo et al. (2022a); Wang et al. (2021, 2023b). However, they struggle to scale with long sequence lengths due to the quadratic complexity of self-attention Vaswani et al. (2017). Various approaches have been proposed to address quadratic complexity, including linear attention Katharopoulos et al. (2020) and methods using sparse attention patterns Guo et al. (2022b); Zaheer et al. (2020); Condevaux and Harispe (2022). LongCoder Guo et al. (2023) uses the sparse attention mechanism. However, efficient transformers have seen limited practical adoption due to a minimal gain in computational cost in practice Yang et al. (2025) and a potential degradation in performance Qin et al. (2022). Dao et al. (2022; 2024) introduced hardware-level optimizations, but the benefits of these optimizations are hardware-specific and do not necessarily apply to newer or different hardware architectures Shah et al. (2024).
@@
Guan et al. (2024) and Wang et al. (2024) enhances the base transformer models for longer contexts using repository-level information retrieval, while Clement et al. (2021); Cheng et al. (2024); Lomshakov et al. (2024); Bi et al. (2024); Liu et al. (2024a); Ma et al. (2024) use static analysis tools. The usage of static analysis tools is complementary to CodeSSM, as the performance of CodeSSM can also be further improved through the use of these tools. Additionally, CodeSSM can be a drop-in replacement as an efficient retriever model.
@@ 7 Conclusions
State-space models (SSMs) have emerged as an efficient alternative to transformers. In this work, we investigated the effectiveness of SSM in code understanding tasks by introducing CodeSSM. Our results show significant potential of CodeSSM in code modeling and understanding. Our findings show that SSMs can be a robust and sample-efficient alternative to transformers, offering the added benefit of length extrapolation.
While the training pipeline and architectural improvements for transformers have evolved extensively over the past decade, the optimal training conditions for SSMs have yet to be thoroughly investigated. This work represents a step toward understanding these conditions and uncovering the advantages of SSMs. Notably, the performance gains achieved by CodeSSM were obtained under suboptimal conditions, indicating the potential for further enhancements. With advancements in training methodologies and architectural designs for SSM-based models, their performance could be significantly improved.
@@
While we have done an extensive study of SSMs for code understanding, our work has some limitations.
  • Single kernel transformation: We have used a simple implementation of S4d Gu et al. (2022a) in the CodeSSM layer. This implementation has only one SSM kernel shared across all input channel. Incorporating multiple kernels in an SSM block can further improve the performance.
  • Short-range dependencies: We have discussed SSMs’ limitations in understanding short-range dependencies with respect to type inference task. However, we do not implement a solution to this problem. Some methods to alleviate this issue have been proposed in literature. We leave the exploration of these methods in the context of code understanding for future work.
@@
We have used publicly existing datasets. CodeSearchNet dataset can contain personally identifiable information (PII). While StarCoder dataset attempts to remove PII, there might still be some PII in the dataset. It is not clear whether our model can leak such data if present in the pre-training dataset, but we acknowledge that such potential risk exist. Further, our work involved training multiple models. While the model and dataset sizes were small, the training can still have environmental impacts.
@@ Appendix C Dataset
Finetuning Datasets. The details about finetuning datasets is presented in Table 6.
@@ Appendix E Pretraining evaluation
In Table 7, we present the pretraining accuracy of BertCoder and multiple CodeSSM variants on a hold-out set from the CodeSearchNet dataset.
@@ Appendix F Theoretical basis for improvements in performance
The plots of magnitude and phase of all layers of the pretrained CodeSSM model for a kernel of length 10 is shown in Figure 4 and Figure 5 respectively. The plot shows forward kernel on the left and backward kernel on the right.
Do note, however, that $\mathbf{A(u)}$ explicitly calculates relation between each pair of input tokens, while $\mathbf{H(\omega)}$ does not. This results in a key limitation of SSMs which is that they fail to encode complex local relations. This failure can result in slight degradation in performance on tasks requiring local information with complex relations among tokens, as evidenced by performance on Type Inference (Table 2).
@@ Appendix G Qualitative Analysis of Failure cases in Type Inference
To predict the correct type, the model must look at the location from where the value of the variable is coming, i.e, the model should encode which tokens influence the value of a given token. Each SSM kernel is a mixture of multiple wavelengths. The influence of one token on another depends on the frequency coefficients at a specific wavelength Hua et al. (2025b). Since our SSM implementation has only one kernel per SSM block, there are only few wavelengths with high coefficients. In Figure 4, it can be observed that the later layers has higher coefficients for lower frequency (longer wavelengths) and hence, focus on long range dependencies. Since, type inference require both long and short range dependency, the model fails at encoding sufficient information. Including multiple kernels in SSM blocks can potentially alleviate this issue.
@@ Appendix H Additional Results
Here are some additional results on Type Inference (Table 8) with Top 100 F1 and on Clone Detection (Table 9) with Precision and Recall.

### v3

@@ 4.3 Token classification task (Type inference)
Column 5 of Table˜2 shows the results on the type inference benchmark. CodeSSM (across all context lengths) matches or exceeds CodeBERT. However, both RoCoder and BertCoder outperform CodeSSM. We provide reasons for the poor performance of CodeSSM in type inference in Section˜5.6 and Appendix˜G.
@@ 4.4 Clone Detection
Column 4 of Table˜2 shows the F1 in the clone detection task. The CodeSSM-sc with a pretraining context length of 256 outperforms both CodeBERT, BertCoder, and RoCoder in terms of F1. Under a similar training setup (2048 context in CSN), removing positional embedding does have a minimal impact on performance, but with the advantage that the maximum input length need not be fixed during pretraining. The decrease in performance after introducing dropout shows that the model does not overfit despite the data set being unbalanced.
@@ 5.1 How sample efficient is CodeSSM?
Figure˜3 shows the MLM training accuracy trajectories of CodeSSM and BertCoder, both trained with a maximum context of 2048. CodeSSM demonstrates rapid initial learning, achieving more than 50% accuracy in only 3,000 training steps (96,000 training samples). BertCoder, on the other hand, exhibits better long-term learning characteristics, outperforming CodeSSM by 4% points after six training epochs (see Appendix˜E). However, CodeSSM outperforms transformers in all downstream tasks except type inference, as discussed in Section˜4. Performance in downstream tasks shows that the code modeling capability observed during pretraining transfers well to multiple downstream tasks.
@@
Our analyses demonstrate that SSMs represent a sample-efficient alternative to transformers for code language modeling and understanding tasks. We attribute the sample efficiency of CodeSSM to the distinct inductive bias of SSM models compared to transformer architectures Wang et al. (2023a). We also present some (informal) theoretical basis for CodeSSM’s performance and the magnitude and phase spectrum of Fourier transform of SSM kernel in Appendix˜F.
@@ 5.2 How well does CodeSSM extrapolate to longer contexts?
We studied transformer models and CodeSSMs for positional embedding and different pretraining context lengths to investigate their length extrapolation capabilities.
@@
Out of two retrieval tasks, RoCoder performs slightly better on NLCodeSearch while CodeSSM outperforms it on SQA. The context length is 512 for the former and 1024 for the latter. Thus, CodeSSM is better at handling unseen input lengths as the context for the downstream task increases. Since SSMs have linear complexity with respect to input length, they are advantageous in terms of both computational efficiency and performance. Moreover, with a longer pretraining context, CodeSSM outperforms RoCoder on NLCodeSearch as well. Therefore, with more computational resources, CodeSSM can be efficiently trained with longer contexts to improve the performance on retrieval tasks. However, it is important to note that longer pretraining contexts do not benefit all downstream tasks; in fact, performance slightly decreases in vulnerability detection and clone detection.
@@
5.3 How fast and memory efficient is CodeSSM?
Table 6: Memory Consumption during Inference (in GB) on various tasks with varying context lengths.
Model NLCodeSearch (512) SQA (1024) SQA (2048)
CodeSSM 9.33 15.84 28.84
RoCoder 9.87 25.36 79.12
The usage of GPU memory during inference is shown in Table˜6. CodeSSM model consumes 5.5% less memory than the transformer model (RoCoder) on the NLCodeSearch task where the context length is 512. The difference in memory consumption between CodeSSM and RoCoder increases to 37.5% on SQA task where the context length is 1024. As the input length increases to 2048, CodeSSM saves 63.55% memory compared to the transformers. These findings are in line with previous works on SSM Le Bronnec et al. (2024).
Table 7: Inference throughput (samples per second) with varying context lengths.
Model NLCodeSearch (512) SQA (1024) SQA (2048)
CodeSSM 36.38 22.50 11.49
RoCoder 38.11 19.76 8.48
Transformers are slightly faster compared to SSM in small input sizes (<=512) but SSMs are faster as the input length increases.
@@ 5.4 Can SSMs be replaced by DFT?
We ask this question because Lee-Thorp et al. (2022) showed that self-attention can be replaced with DFT in transformers with minimal reduction in NLP tasks. We also observed that in all tasks RoCoder performed significantly better than BertCoder. Hua et al. (2025b) showed that RoPE also performs token-level (nonuniform) DFT. Additionally, SSMs also implicitly perform DFT of input and kernel during convolution operation.
To answer this question, we performed an ablation study by replacing the SSM transformation in CodeSSM with DFT and creating a unidirectional variant of CodeSSM. The results are presented in Figure˜4 (and in Table˜10). We observe that both variants perform significantly worse compared to CodeSSM. However, it is important to note that CodeF performs better than CodeSSM-uni, i.e., the advantages of SSM over DFT on code modeling are only realized with bidirectional SSMs.
@@ 5.5 Is CodeSSM robust to overfitting on small datasets?
We study the robustness of the CodeSSM model in two ways. First, we study the impact of adding dropout to the CodeSSM layer. Across all tasks, including the complexity prediction task, which has only 3.6k samples, the performance of the model decreases when dropout is added. This performance degradation suggests that CodeSSM does not overfit during training regardless of the size of the finetuning dataset. Second, we evaluated the model on the adversarial NLCodeSearch task, by which informative tokens such as function and variable names have been replaced with noninformative generic tokens. The CodeSSM model performs very well in this task even with minimal training.
@@ 5.6 When does CodeSSM fail and why?
As discussed in Section˜5.1, transformers can outperform SSMs as the number of samples increases. This phenomenon can hinder the performance of SSMs on some tasks such as Type Inference. This benchmark consists of over 9M type annotations, which is significantly larger than the samples of other tasks. The total number of predictions (hence, feedback to the model through loss) is even larger since there is one prediction for every token. However, when the number of samples is large and the context length of finetuning datasets is long, SSMs can surpass transformers in such scenarios (as seen in the SQA benchmark).
SSMs can effectively capture long-range dependencies, but they do not model local dependencies (Zuo et al., 2024) very well. The type inference benchmark requires the model to understand both short- and long-range dependencies. For example, in the code sample x = y+1; z = func(x), the type of x can be inferred from the local context (variable y), while the type of z can only be inferred from the return type of the function func, which requires capturing long-range dependencies. We also observe that this problem becomes more challenging when the local context is complicated. We provide a qualitative example of the failure of CodeSSM in the type inference task in Appendix˜G. We can potentially improve CodeSSM’s performance on this task by making the step size of SSMs (see Appendix˜A) data dependent Qi et al. (2024) or with multiple SSM kernels (Appendix˜G), which can improve its understanding of short-range dependencies.
5.7 Expressivity gaps in SSMs and their fixes
Some recent theoretical works document the expressivity limitations of SSMs. In particular, Merrill et al. (2024) shows that SSMs cannot express some important problems that require state tracking. This class of problems includes evaluating Python code. However, it should be noted that transformers have also been shown to not express this class of problems while a single layer RNN can Merrill and Sabharwal (2023). However, the expressive power of SSMs can be improved through different architectural improvements. For example, Merrill et al. (2024) showed that making the input of the transition matrix independent or adding nonlinearities can extend the expressive power of SSM. Similarly, Nishikawa and Suzuki (2025) showed that when SSMs are preceded and followed by feed-forward layers, they have the same dynamic token abilities as transformers.
Another limitation is the requirement of specific reparameterization. Agarwal et al. (2024) proposed a new SSM formulation by augmenting spectral filters with negative eigenvalues, which removed the requirement of reparameterization. The SSM variant used in CodeSSM is S4D. S4D also has negative eigenvalues but still requires specific reparameterization. The best way to parametrize SSM and their practical benefits remain an open research question Wang and Li (2024).
@@ 6 Related Work
Transformer models have been used in code intelligence tasks Feng et al. (2020); Guo et al. (2022a); Wang et al. (2021, 2023b). However, they struggle to scale with long sequence lengths due to the quadratic complexity of self-attention Vaswani et al. (2017). Various approaches have been proposed to address quadratic complexity, including linear attention Katharopoulos et al. (2020) and methods using sparse attention patterns Guo et al. (2022b); Zaheer et al. (2020); Condevaux and Harispe (2022). LongCoder Guo et al. (2023) uses the sparse attention mechanism.
However, efficient transformers have seen limited practical adoption due to a minimal gain in computational cost in practice Yang et al. (2025) and a potential degradation in performance Qin et al. (2022). Dao et al. (2022) and Dao (2024) introduced hardware-level optimizations, but the benefits of these optimizations are hardware-specific and do not necessarily apply to newer or different hardware architectures Shah et al. (2024).
@@
Guan et al. (2024) and Wang et al. (2024) improve base transformer models for longer contexts using repository-level information retrieval, while Clement et al. (2021); Cheng et al. (2024); Lomshakov et al. (2024); Bi et al. (2024); Liu et al. (2024a); Ma et al. (2024) use static analysis tools for repository-level tasks. The usage of static analysis tools is complementary to CodeSSM, as the performance of CodeSSM can also be further improved by using these tools. Additionally, CodeSSM can be a drop-in replacement as an efficient retriever model.
@@ 7 Conclusions
State-space models (SSMs) have emerged as an efficient alternative to transformers. In this work, we investigated the effectiveness of SSM in code understanding tasks by introducing CodeSSM. Our results show a significant potential for CodeSSM in code modeling and understanding. Our findings show that SSMs can be a robust and sample-efficient alternative to transformers, offering the added benefit of length extrapolation.
Although the training pipeline and architectural improvements for transformers have evolved extensively in the past decade, the optimal training conditions for SSMs have yet to be thoroughly investigated. This work represents a step toward understanding these conditions and uncovering the advantages of SSMs. In particular, the performance gains achieved by CodeSSM were obtained under suboptimal conditions, indicating the potential for further enhancements. With advancements in training methodologies and architectural designs for SSM-based models, their performance could be significantly improved.
@@
Although we have done an extensive study of SSMs for code understanding, our work has some limitations.
  • Single kernel transformation: We have used a simple implementation of S4d Gu et al. (2022a) in the CodeSSM layer. This implementation has only one SSM kernel shared across all input channels. Incorporating multiple kernels into an SSM block can further improve performance.
  • Short-range dependencies: We have discussed SSMs’ limitations in understanding short-range dependencies in the context of the type inference task. However, we do not provide a solution to this problem. While some methods have been proposed to alleviate this issue in the literature, we leave the their exploration in the context of code understanding for future work.
@@
We have used publicly existing datasets. CodeSearchNet dataset can contain personally identifiable information (PII). While the StarCoder dataset attempts to remove PII, there might still be some PII in the dataset. It is not clear whether our model can leak such data if present in the pre-training dataset, but we acknowledge that such potential risk exists. In addition, our work involved training multiple models. Although the size of the model and the data set was small, training can still have environmental impacts.
@@ Appendix C Dataset
Finetuning Datasets. The details about finetuning datasets is presented in Table˜8.
@@ Appendix D Other Training Setup and Details
Table 9: Hyperparameters for CodeSSM, BertCoder, RoCoder and Zamba models
Hyperparameters CodeSSM BertCoder RoCoder Zamba
Number of Layers 12 12 12 14
Hidden Dimension 1024 1024 1032 1024
Positional Embedding No Yes Yes No
Pretraining context length 256 2048 256 256
@@ Appendix E Pretraining evaluation
In Table˜10, we present the pretraining accuracy of BertCoder and multiple CodeSSM variants on a hold-out set from the CodeSearchNet dataset.
@@ Appendix F Theoretical basis for improvements in performance
The plots of magnitude and phase of all layers of the pretrained CodeSSM model for a kernel of length 10 is shown in Figure˜5 and Figure˜6 respectively. The plot shows forward kernel on the left and backward kernel on the right.
Do note, however, that $\mathbf{A(u)}$ explicitly calculates relation between each pair of input tokens, while $\mathbf{H(\omega)}$ does not. This results in a key limitation of SSMs which is that they fail to encode complex local relations. This failure can result in slight degradation in performance on tasks requiring local information with complex relations among tokens, as evidenced by performance on Type Inference (Table˜2).
@@ Appendix G Qualitative Analysis of Failure cases in Type Inference
To predict the correct type, the model must look at the location from where the value of the variable is coming, i.e, the model should encode which tokens influence the value of a given token. Each SSM kernel is a mixture of multiple wavelengths. The influence of one token on another depends on the frequency coefficients at a specific wavelength Hua et al. (2025b). Since our SSM implementation has only one kernel per SSM block, there are only few wavelengths with high coefficients. In Figure˜5, it can be observed that the later layers has higher coefficients for lower frequency (longer wavelengths) and hence, focus on long range dependencies. Since, type inference require both long and short range dependency, the model fails at encoding sufficient information. Including multiple kernels in SSM blocks can potentially alleviate this issue.
@@ Appendix H Additional Results
Here are some additional results on Type Inference (Table˜11) with Top 100 F1 and on Clone Detection (Table˜12) with Precision and Recall.